# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 2 Preliminaries
[…]
@@
[8, Section 13] The search operation takes $O(\mathrm{log}\;n)$ time on a red-black tree, where $n$ is the size of the tree.
@@
[8, Section 13] The time-complexity of inserting $n\in{\mathbb{N}}$ elements into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\;n)$.
@@
A binary sequence is a string of ones and zeros, $BIN$ stands for the set of binary sequences. For an integer $\ell\in{\mathbb{N}}$, we write $BIN_{\ell}$ to denote the set of binary sequences of length $\ell$. We define the binary sequence $0^{k}$, $k\in{\mathbb{N}}$ by recursion on $k$. Let $0^{0}=\varepsilon$. For each $k\in{\mathbb{N}}$, let $0^{k+1}=00^{k}$. Let $0\leq k\leq\ell$ and $\alpha\in BIN_{k}$. We pad the binary sequence $\alpha$ with leading zeros to length $\ell$ by producing the binary sequence $0^{\ell-k}\alpha$. We now recall the definition of the lexicographic order on $BIN$. For $\alpha,\beta\in BIN$, we say that $\alpha$ is lexicographically less than $\beta$, denoted by $\alpha<_{lex}\beta$ if either $\alpha$ is a proper prefix of $\beta$, or there exist $\gamma_{1},\gamma_{2},\gamma_{3}\in BIN$ such that $\alpha=\gamma_{1}0\gamma_{2}$ and $\beta=\gamma_{1}1\gamma_{3}$.
[…]
  • Noetherian if there exists no infinite sequence of elements $a_{1},a_{2},a_{3},\ldots$ in $A$ such that $a_{1}{\rightarrow}a_{2}{\rightarrow}a_{3}{\rightarrow}\ldots$,
@@
  • confluent if for any elements $a_{1},a_{2},a_{3}$ in $A$, whenever $a_{1}{\rightarrow}^{*}a_{2}$ and $a_{1}{\rightarrow}^{*}a_{3}$, there exists an element $a_{4}$ in $A$ such that $a_{2}{\rightarrow}^{*}a_{4}$ and $a_{3}{\rightarrow}^{*}a_{4}$,
  • convergent if it is Noetherian and confluent.
[…]
Proposition 2.3.
[…]
If $a\in A$ has a unique normalform, then we denote the latter by $a{\downarrow}$.
@@
Let $\rho$ be an equivalence relation on $A$. Then for every $a\in A$, we denote by $a/_{\rho}$ the $\rho$ equivalence class containing $a$, i.e. $a/_{\rho}=\{\,b\mid a\rho b\,\}$. The classes associated with an equivalence relation $\rho$ are called equivalence classes. For each $B\subseteq A$, let $B/_{\rho}=\{\,b/_{\rho}\mid b\in B\,\}$.
@@
The empty product $\Pi_{i=1}^{0}n_{i}=1$ over ${\mathbb{N}}$ is the result of multiplying no factors. It is by convention equal to $1$. The ceiling function maps a real number $x$ to the least integer greater than or equal to $x$, denoted $\lceil x\rceil$.
Ground terms
@@
A signature $\Sigma$ is a finite set of symbols in which every element $\sigma$ has an arity, $\mathrm{arity}(\sigma)\in{\mathbb{N}}$. For each integer $m\in{\mathbb{N}}$, $\Sigma_{m}$ denotes the elements of $\Sigma$ which have arity $m$. We lose no generality assuming that $\Sigma_{0}\not=\emptyset$, see the Introduction. If $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, then $\Sigma$ is a unary signature. $mar_{\Sigma}$ denotes the maximal arity of the symbols in $\Sigma$.
@@
    $\sigma(t_{1},\ldots,t_{m})\in U$ whenever $\sigma\in\Sigma_{m}$ with $m\in{\mathbb{N}}$ and $t_{1},\ldots,t_{m}\in U$.
@@
  • If $t=\sigma(t_{1},\ldots,t_{m})$ with $\sigma\in\Sigma_{m},\ m>0,$ then $\mathrm{POS}(t)=\{\,\varepsilon\,\}\cup\{\,i\alpha\mid 1\leq i\leq m,\alpha\in\mathrm{POS}(t_{i})\,\}$.
The size of a ground term $t\in T_{\Sigma}$ is $size(t)=|POS(t)|$. For a finite set $Z\subseteq T_{\Sigma}$ of ground terms, $size(Z)$ is the number of occurrences of symbols of $\Sigma$ in $Z$, i.e., $size(Z)=\sum_{t\in Z}size(t)$. For any ground term $t\in T_{\Sigma}$ and position $\alpha\in\mathrm{POS}(t)$, we introduce the subterm $t|_{\alpha}\in T_{\Sigma}$ of $t$ at $\alpha$ as follows:
@@
A context $\delta$ over $\Sigma$ can be viewed as an incomplete term, which may contain several empty places or holes. Formally, a context over $\Sigma$ is a ground term containing zero, one or more occurrences of a special constant symbol $\diamondsuit\not\in\Sigma$ denoting holes, i.e., a ground term over the extended signature $\Sigma\cup\{\,\diamondsuit\,\}$. If $\delta$ contains exactly $n\in{\mathbb{N}}$ holes, then we say that $\delta$ is an $n$-context. The set of all $n$-contexts over $\Sigma$ is denoted by $CON_{\Sigma,n}$. For example, let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\$\,\}$, and $\Sigma_{2}=\{\,f,g\,\}$. Then $g(f(\diamondsuit,\#),g(\diamondsuit,\diamondsuit))\in CON_{\Sigma,3}$. Let $\delta$ be an $n$-context over $\Sigma$, and let $t_{1},\ldots,t_{n}\in T_{\Sigma}$ be ground terms. Then $\delta[t_{1},\ldots,t_{n}]$ denotes the result of replacing the holes of $\delta$ from left to right by $t_{1},\ldots,t_{n}$. For example,
$g(f(\diamondsuit,\#),f(\diamondsuit,\diamondsuit))[g(\$,\$),\#,\$]=g(f(g(\$,\$),\#),f(\#,\$))$.
Let $\delta$ be a $1$-context over $\Sigma$, and let $\alpha\in POS(\delta)$ be such that $t|_{\alpha}=\diamondsuit$. Then we call $\alpha$ the address of $\diamondsuit$ in $\delta$, and $addr(\delta)$ stands for $\alpha$. Let $\delta$ be a $1$-context over $\Sigma$ such that $\delta\neq\diamondsuit$, and let $c\in\Sigma_{0}$ be a constant. Then we say that the ground term $\delta[c]$ is a proper extension of the constant $c\in\Sigma_{0}$.
@@
Let $\Sigma$ be a signature, and $R$ be a finite subset of $T_{\Sigma}\times T_{\Sigma}$. The size $size(R)$ of $R$ is the number of occurrences of symbols of $\Sigma$ in $R$, ie., $size(R)=\sum_{(l,r)\in R}size(l)+size(r)$. Let $\Sigma\langle R\rangle$ denote the signature of those symbols which occur in $R$. For any ground terms $s,t\in T_{\Sigma}$, we say that $s$ rewrites to $t$, and denote this by $s\mathop{\rightarrow}_{R}t$ if there exist an ordered pair $(l,r)$ in $R$, a context $\delta\in CON_{\Sigma,1}$ such that $s=\delta[l]$ and $t=\delta[r]$. We fully specify such a rewrite step by $s\rightarrow_{[\alpha,(l,r)]}t$, where $\alpha=addr(\delta)$. Furthermore, $s\leftrightarrow_{[\alpha,(l,r)]}t$ stands for ($s\rightarrow_{[\alpha,(l,r)]}t$ or $t\rightarrow_{[\alpha,(l,r)]}s$). When we focus on the relations $\mathop{\rightarrow}_{R}$ and $\mathop{\rightarrow}_{R}^{*}$, we call $R$ a ground term rewrite system (GTRS for short), and we call an ordered pair $(l,r)$ a rewrite rule, and denote it as $l\dot{\rightarrow}r$. Moreover, we say that $l$ is the left-hand side and $r$ is the right-hand side of the rewrite rule $l\dot{\rightarrow}r$. When we focus on the symmetric relations $\mathop{\leftrightarrow}_{R}$ and $\mathop{\leftrightarrow}^{*}_{R}$, we call $R$ a ground term equation system (GTES for short), and we denote $(l,r)$ as $l\doteq r$, and call it an equation. Moreover, we denote a GTES as $E$ rather than $R$ [3].
Statement 2.4.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. We can construct the GTES $E\cup F$ in $O(n\,\mathrm{log}\,n)$ time.
@@
In $O(n\,\mathrm{log}\,n)$ time we mergesort the GTES $E$ and the GTES $F$ [10, Section 2.2]. Then in $O(n)$ time, scan through each of the sorted relations to produce the result. When the concurrent scan of both GTESs reveals the same equation in both relations, only one of the equations is retained. ∎
@@
Statement 2.6.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=size(E)+size(F)$. We can count the elements of $\Sigma\langle E,F\rangle$ and compute the Boolean value $FLAG\langle E,F\rangle$ in $O(n)$ time.
@@
is the set of subterms occurring $E$.
Statement 2.7.
1. Let $Z\subseteq T_{\Sigma}$ be a finite set of ground terms over a signature $\Sigma$. Then $|sub(Z)|\leq size(Z)$.
2. Let $E$ be a GTES over a signature $\Sigma$. Then $|ST\langle E\rangle|\leq size(E)$.
Proof.
By definition. ∎
Let $\Sigma$ be a signature. A $\Sigma$-algebra ${\cal A}$ consists of a carrier set $A$, and a mapping that associates with each symbol $\sigma\in\Sigma_{m}$ a mapping $\sigma^{\cal A}:A^{m}\rightarrow A$ for every $m\in{\mathbb{N}}$ [9]. The $\Sigma$-ground-term algebra ${\cal T}(\Sigma)$ has $T_{\Sigma}$ as carrier set, and interprets the symbols $\sigma\in\Sigma_{m}$ for every $m\in{\mathbb{N}}$ as follows:
$\sigma^{{\cal T}(\Sigma)}:T_{\Sigma}^{m}\rightarrow T_{\Sigma}:(t_{1},\ldots,t_{m})\mapsto\sigma(t_{1},\ldots,t_{m})$.
Let $E$ be a GTES over $\Sigma$. The congruence relation generated by $E$ on the $\Sigma$-ground term algebra ${\cal T}(\Sigma)$ is $\mathop{\leftrightarrow}^{*}_{E}$. It is well known that $\mathop{\leftrightarrow}^{*}_{E}$ is the least congruence on ${\cal T}(\Sigma)$ containing $E$. Let $R$ be a GTRS $R$ over $\Sigma$. We say that GTRS $R$ is equivalent to the GTES $E$ if $\mathop{\leftrightarrow}^{*}_{R}=\mathop{\leftrightarrow}^{*}_{E}$. We say that GTRS $R$ reaches a ground term $s\in T_{\Sigma}$ from a ground term $t\in T_{\Sigma}$ if $t\mathop{\rightarrow}_{R}^{*}s$ holds. We say that GTRS $R$ is Noetherian, (confluent, etc.) if $\mathop{\rightarrow}_{R}$ is Noetherian (confluent, etc.). A term $t\in T_{\Sigma}$ is irreducible with respect to $R$ if it is irreducible with respect to $\mathop{\rightarrow}_{R}$. If a term $t\in T_{\Sigma}$ has a unique $\mathop{\rightarrow}_{R}$ normal form, then the latter is denoted by $t{\downarrow}_{R}$. A GTRS $R$ is reduced if for every rewrite rule $l\dot{\rightarrow}r$ in $R$, $l$ is irreducible with respect to $R-\{\,l\dot{\rightarrow}r\,\}$ and $r$ is irreducible for $R$. We recall the following important result.
Proposition 2.8.
[3] Every reduced GTRS $R$ is convergent.
@@
Let $\Sigma$ be a signature, a bottom-up tree automaton (bta for short) over $\Sigma$ is a quadruple ${\cal A}=(\Sigma,A,R,A_{f})$, where $A$ is a finite set of states of arity 00, $\Sigma\cap A=\emptyset$, $A_{f}(\subseteq A)$ is the set of final states, $R$ is a finite set of rewrite rules of the following form:
$\sigma(a_{1},\ldots,a_{m})\dot{\rightarrow}a$ with $m\in{\mathbb{N}}$, $\sigma\in\Sigma_{m}$, $a_{1},\ldots,a_{m},a\in A$.
We consider $R$ as a GTRS over $\Sigma\cup A$. The size $size({\cal A})$ of $\cal A$ is $|A|+size(R)$. The tree language recognized by $\cal A$ is $L({\cal A})=\{\,t\in T_{\Sigma}\mid(\exists a\in A_{f})\;t\mathop{\rightarrow}_{R}^{*}a\}$. We say that ${\cal A}=(\Sigma,A,R,A_{f})$ is a deterministic bta if for any $\sigma\in\Sigma_{m}$, $m\in{\mathbb{N}}$, $a_{1},\ldots,a_{m}\in A$, there is at most one rule with left-hand side $\sigma(a_{1},\ldots,a_{m})$ in $R$.
@@
[13] Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let $W=ST(E\cup F)$. For every $p\in W$, we can construct in $O(n\,\mathrm{log}\,n)$ time the deterministic bottom-up tree automaton ${\cal A}=(\Sigma,C\langle E,F\rangle,R\langle E,F\rangle,\{\,p/_{\Theta\langle E,F\rangle}\,\})$, and ${\cal A}$ recognizes the tree language $p/_{\mathop{\leftrightarrow}^{*}_{E}}$.
@@
[5] Given two btas ${\cal A}=(\Sigma,A,R_{\cal A},A_{f})$ and ${\cal B}=(\Sigma,B,R_{\cal B},B_{f})$, we can decide whether $L({\cal A})\subseteq L({\cal B})$ in $O(size({\cal A})\cdot size({\cal B}))$ time.
@@
A directed pseudograph without parallel arcs (dpwpa for short) $G$ is an ordered pair $(V,A)$, where $V$ is a finite set, and $A\subseteq V\times V$ is a relation on $V$. The set $V$ is called the vertex set of $G$, and its elements are called vertices. The set $A$ is called the directed arc set of $G$, and its elements are called directed arcs. A walk in $G$ is an alternating sequence $w=x_{1}\,(x_{1},x_{2})\,x_{2}\,(x_{2},x_{3})\,x_{3}\,\ldots\,x_{k-1}\,(x_{k-1},x_{k})\,x_{k}$, $k\geq 1$, of vertices $x_{i}$ and arcs $(x_{j},x_{j+1})$. We say that $w$ is from vertex $x_{1}$ to vertex $x_{k}$. The length of $w$ is the number $k-1$ of its arcs. A trail is a walk $w$ in which all arcs are distinct. If the vertices of a trail $w$ are distinct, then $w$ is a path. If the vertices $x_{1},x_{2},\ldots,x_{k-1}$ of a trail $w$ are distinct, $k\geq 2$, and $x_{1}=x_{k}$, then $w$ is a cycle. We say that a vertex $x$ is reachable from a vertex $y$ if there exists a walk $w$ from vertex $y$ to vertex $x$. In particular, a vertex is reachable from itself. We say that a vertex $x$ is positive-step reachable from a vertex $y$ if there exists a walk $w$ of positive length from vertex $y$ to vertex $x$. Observe that for any vertices $x$ and $y$, the vertex $x$ is positive-step reachable from the vertex $y$ if and only if there exists a path of positive length or a cycle from the vertex $x$ to the vertex $y$.

### v2

@@ 2 Preliminaries
In algorithmic analysis we use the RAM (random access machine) model [11, Section 2.2], [12, Section 2.1]. For this model of computation, each simple operation (+,,,=,if,call)(+,\,*,\,–,\,=,\,if,\,call) takes exactly one time step. An array is a collection of items of same data type stored at contiguous memory locations. For the concept of a red-black tree, see [13, Section 13].
@@
[13, Section 13] The search operation takes $O(\mathrm{log}\,n)$ time on a red-black tree, where $n$ is the size of the tree.
@@
[13, Section 13] The time-complexity of inserting $n\in{\mathbb{N}}$ elements into an empty red-black tree, not allowing duplicate values in the resulting red-black tree, is $O(n\,\mathrm{log}\,n)$.
@@
A binary sequence is a string of ones and zeros, $BIN$ stands for the set of binary sequences. For an integer $\ell\in{\mathbb{N}}$, we write $BIN_{\ell}$ to denote the set of binary sequences of length $\ell$. We define the binary sequence $0^{k}$, $k\in{\mathbb{N}}$ by recursion on $k$. Let $0^{0}=\varepsilon$. For each $k\in{\mathbb{N}}$, let $0^{k+1}=00^{k}$. Let $0\leq k\leq\ell$ and $\alpha\in BIN_{k}$. We pad the binary sequence $\alpha$ with leading zeros to length $\ell$ by producing the binary sequence $0^{\ell-k}\alpha$.
We now recall the definition of the lexicographic order on $BIN$ [14, Definition 6.4.1.]. Given two distinct binary strings, to determine whether one precedes the other in the lexicographic order we compare the numbers in the words. There are two possibilities:
  1. 1.
    The words have different lengths, and each number in the shorter word is identical to the corresponding number in the longer word.
  2. 2.
    The words have the same or different lengths, and at some position, the numbers in the words differ.
If Condition 1 holds, then the shorter word precedes the longer. If Condition 2 holds, we locate the leftmost position $\ell$ at which the numbers differ. The order of the words is determined by the order of the numbers at position $\ell$.
Definition 2.3 ([14, Definition 6.4.1.]).
Let $\alpha=i_{1}i_{2}\ldots i_{m}$ and $\beta=j_{1}j_{2}\ldots j_{n}$ be binary strings. We say that $\alpha$ is lexicographically less than $\beta$ and write $\alpha<_{lex}\beta$ if
  1. 1.
    $m<n$ and $i_{k}=j_{k}$ for every $k=1,2,\ldots,m$ or
  2. 2.
    $i_{k}\neq j_{k}$ for some $k$ and for the smallest such $k$, we have $i_{k}<j_{k}$.
Sets and Relations [15]
For any sets $A$ and $B$, $A\subseteq B$ means that $A$ is a subset of $B$. Proper inclusion is denoted by $A\subset B$. The union of two sets $A$ and $B$ is denoted by $A\cup B$. The cardinality of a set $A$ is denoted $|A|$. A relation on a set $A$ is a subset $\rightarrow$} of $A\times A$ [15]. We write ${b}{\rightarrow}{c}$ for $({b},{c})\in\,{\rightarrow}$. We denote by $\rightarrow$$^{*}$ the reflexive transitive closure of $\rightarrow$}, by $\leftrightarrow^{+}$ the transitive symmetric closure of $\rightarrow$}, and by $\leftrightarrow^{*}$ the reflexive transitive symmetric closure of $\rightarrow$}. Note that $\leftrightarrow^{*}$ is an equivalence relation on $A$.
@@
Let $\rightarrow$} be a relation on a set $A$. An element ${b}\in A$ is in normal form (irreducible for $\rightarrow$}) if there exists no ${c}\in A$ such that ${b}{\rightarrow}{c}$. We say that ${c}\in A$ is a normal form of ${b}\in A$ if ${b}{\rightarrow}^{*}{c}$ and c{c} is in normal form.
A relation $\rightarrow$} on a set $A$ is
  • 1.
    normalizing if every element has a normal form,
  • 2.
    terminating if there exists no infinite sequence of elements ${b}_{1},{b}_{2},{b}_{3},\ldots$ in $A$ such that ${b}_{1}{\rightarrow}{b}_{2}{\rightarrow}{b}_{3}{\rightarrow}\ldots$,
  • 3.
    confluent if for any elements ${b}_{1},{b}_{2},{b}_{3}$ in $A$, whenever ${b}_{1}{\rightarrow}^{*}{b}_{2}$ and ${b}_{1}{\rightarrow}^{*}{b}_{3}$, there exists an element ${b}_{4}$ in $A$ such that ${b}_{2}{\rightarrow}^{*}{b}_{4}$ and ${b}_{3}{\rightarrow}^{*}{b}_{4}$,
  • 4.
    convergent if it is terminating and confluent.
Any terminating relation is normalizing. Hence we have the following proposition.
Proposition 2.4 ([15, Lemma 2.1.8]).
For any convergent relation $\rightarrow$} on a set $A$, every ${b}\in A$ has a unique normal form, called the normal form of b{b}.
If ${b}\in A$ has a unique normalform, then we denote the latter by ${b}{\downarrow}$.
@@
Let $\rho$ be an equivalence relation on $A$. Then for every ${b}\in A$, we denote by ${b}/_{\rho}$ the $\rho$ equivalence class containing b{b}, i.e. ${b}/_{\rho}=\{\,{c}\mid{b}\rho{c}\,\}$. For each $B\subseteq A$, let $B/_{\rho}=\{\,b/_{\rho}\mid b\in B\,\}$.
@@
The empty product $\Pi_{i=1}^{0}n_{i}=1$ over $\mathbb{N}$} is the result of multiplying no factors. The ceiling function maps a real number $x$ to the least integer greater than or equal to $x$, denoted $\lceil x\rceil$.
Ground terms [15]
@@
A signature $\Sigma$ is a finite set of symbols in which every element $\sigma$ is associated with a non-negative integer, the arity of $\sigma$ denoted by $\mathrm{arity}(\sigma)$. For each integer $k\in{\mathbb{N}}$, $\Sigma_{k}$ denotes the elements of $\Sigma$ which have arity $k$. We lose no generality assuming that $\Sigma_{0}\not$$\emptyset$, see the Introduction. If $\Sigma=\Sigma_{0}\cup\Sigma_{1}$, then $\Sigma$ is a unary signature. $mar_{\Sigma}$ denotes the maximal arity of the symbols in $\Sigma$.
@@
    $\sigma(t_{1},\ldots,t_{m})\in U$ whenever $\sigma\in\Sigma_{m}$ with $m\geq 1$ and $t_{1},\ldots,t_{m}\in U$.
@@
  • 2.
    If $t=\sigma(t_{1},\ldots,t_{m})$ with $\sigma\in\Sigma_{m}$, $m\geq 1$, then $\mathrm{POS}(t)=\{\,\varepsilon\,\}\cup\{\,i\alpha\mid 1\leq i\leq m,\alpha\in\mathrm{POS}(t_{i})\,\}$.
The size of a ground term $t\in T_{\Sigma}$ is $\mathrm{size}(t)=|POS(t)|$. For a finite set $Z\subseteq T_{\Sigma}$ of ground terms, $\mathrm{size}(Z)$ is the number of occurrences of symbols of $\Sigma$ in $Z$, i.e., $\mathrm{size}(Z)=\sum_{t\in Z}\mathrm{size}(t)$. For any ground term $t\in T_{\Sigma}$ and position $\alpha\in\mathrm{POS}(t)$, we introduce the subterm $t|_{\alpha}\in T_{\Sigma}$ of $t$ at $\alpha$ as follows:
  • 1.
@@
A context $\delta$ over $\Sigma$ can be viewed as an incomplete term, which may contain several empty places or holes. Formally, a context over $\Sigma$ is a ground term containing zero, one or more occurrences of a special constant symbol $\diamondsuit\not\in\Sigma$ denoting holes, i.e., a ground term over the extended signature $\Sigma\cup\{\,\diamondsuit\,\}$. If $\delta$ contains exactly $n\in{\mathbb{N}}$ holes, then we say that $\delta$ is an $n$-context. The set of all $n$-contexts over $\Sigma$ is denoted by $CON_{\Sigma,n}$. For example, let $\Sigma=\Sigma_{0}\cup\Sigma_{2}$, $\Sigma_{0}=\{\,\#,\mathdollar\,\}$, and $\Sigma_{2}=\{\,f,g\,\}$. Then $g(f(\diamondsuit,\#),g(\diamondsuit,\diamondsuit))\in CON_{\Sigma,3}$. Let $\delta$ be an $n$-context over $\Sigma$, and let $t_{1},\ldots,t_{n}\in T_{\Sigma}$ be ground terms. Then $\delta[t_{1},\ldots,t_{n}]$ denotes the result of replacing the holes of $\delta$ from left to right by $t_{1},\ldots,t_{n}$. For example,
$)g(f(\diamondsuit,\#),f(\diamondsuit,\diamondsuit))[g(\mathdollar,\mathdollar),\#,\mathdollar]=g(f(g(\mathdollar,\mathdollar),\#),f(\#,\mathdollar$)).
Let $\delta$ be a $1$-context over $\Sigma$, and let $\alpha\in POS(\delta)$ be such that $t|_{\alpha}=\diamondsuit$. Then we call $\alpha$ the address of $\diamondsuit$ in $\delta$, and $addr(\delta)$ stands for $\alpha$. Let $\delta$ be a $1$-context over $\Sigma$ such that $\delta\neq\diamondsuit$, and let ${d}\in\Sigma_{0}$ be a constant. Then we say that the ground term $\delta[{d}]$ is a proper extension of the constant ${d}\in\Sigma_{0}$.
@@
Let $\Sigma$ be a signature, and $R$ be a finite subset of $T_{\Sigma}\times T_{\Sigma}$. The size of $R$ is denoted by $\mathrm{size}(R)$, and is the number of occurrences of symbols of $\Sigma$ in $R$, ie., $\mathrm{size}(R)=\sum_{(l,r)\in R}\mathrm{size}(l)+\mathrm{size}(r)$. Let $\Sigma\langle R\rangle$ denote the signature of those symbols which occur in $R$. For any ground terms $s,t\in T_{\Sigma}$, we say that $s$ rewrites to $t$, and denote this by $s\mathop{\rightarrow}_{R}t$ if there exist an ordered pair $(l,r)$ in $R$, a context $\delta\in CON_{\Sigma,1}$ such that $s=\delta[l]$ and $t=\delta[r]$. We fully specify such a rewrite step by $s\rightarrow_{[\alpha,(l,r)]}t$, where $\alpha=addr(\delta)$. Furthermore, $s\leftrightarrow_{[\alpha,(l,r)]}t$ stands for ($s\rightarrow_{[\alpha,(l,r)]}t$ or $t\rightarrow_{[\alpha,(l,r)]}s$). When we focus on the relations $\mathop{\rightarrow}_{R}$ and $\mathop{\rightarrow}_{R}^{*}$, we call $R$ a ground term rewrite system (GTRS for short), and we call an ordered pair $(l,r)$ a rewrite rule, and denote it as $l\dot{\rightarrow}r$. Moreover, we say that $l$ is the left-hand side and $r$ is the right-hand side of the rewrite rule $l\dot{\rightarrow}r$. When we focus on the symmetric relations $\mathop{\leftrightarrow}_{R}$ and $\mathop{\leftrightarrow}^{*}_{R}$, we call $R$ a ground term equation system (GTES for short), and we denote $(l,r)$ as $l\doteq r$, and call it an equation. Moreover, we denote a GTES as $E$ rather than $R$ [4]. Let $E$ and $F$ be GTESs over a signature $\Sigma$. Then we define the GTES $E\cup F$ over $\Sigma$ as the union of the sets $E$ and $F$.
Statement 2.5.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. We can construct the GTES $E\cup F$ over $\Sigma$ in $O(n\,\mathrm{log}\,n)$ time.
@@
In $O(n\,\mathrm{log}\,n)$ time we mergesort the GTES $E$ and the GTES $F$ [16, Section 2.2]. Then in $O(n)$ time, we concurrently scan through both of the sorted relations to produce the result. When the concurrent scan of both GTESs reveals the same equation in both relations, only one of the equations is retained. ∎
@@
Statement 2.7.
Let $E$ and $F$ be GTESs over a signature $\Sigma$, and $n=\mathrm{size}(E)+\mathrm{size}(F)$. We can count the elements of $\Sigma\langle E,F\rangle$ and compute the Boolean value $FLAG\langle E,F\rangle$ in $O(n)$ time.
@@
is the set of subterms occurring $E$. The below claim is immediate.
Statement 2.8.
1. Let $Z\subseteq T_{\Sigma}$ be a finite set of ground terms over a signature $\Sigma$. Then $|sub(Z)|\leq\mathrm{size}(Z)$.
2. Let $E$ be a GTES over a signature $\Sigma$. Then $|ST\langle E\rangle|\leq\mathrm{size}(E)$.
Let $\Sigma$ be a signature. A $\Sigma$-algebra $\cal A$} consists of a carrier set $A$, and a mapping that associates with each symbol $\sigma\in\Sigma_{k}$ a mapping $\sigma^{\cal A}:A^{k}\rightarrow A$ for every $k\in{\mathbb{N}}$ [15, Definition 3.2.1]. The $\Sigma$-ground-term algebra ${\cal T}(\Sigma)$ has $T_{\Sigma}$ as carrier set, and interprets the symbols $\sigma\in\Sigma_{k}$ for every $k\in{\mathbb{N}}$ as follows:
$\sigma^{{\cal T}(\Sigma)}:T_{\Sigma}^{k}\rightarrow T_{\Sigma}:(t_{1},\ldots,t_{k})\mapsto\sigma(t_{1},\ldots,t_{k})$.
Let $E$ be a GTES over $\Sigma$. The congruence relation generated by $E$ on the $\Sigma$-ground term algebra ${\cal T}(\Sigma)$ is $\mathop{\leftrightarrow}^{*}_{E}$. It is well known that $\mathop{\leftrightarrow}^{*}_{E}$ is the least congruence on ${\cal T}(\Sigma)$ containing $E$. Let $R$ be a GTRS over $\Sigma$. We say that GTRS $R$ is equivalent to the GTES $E$ if $\mathop{\leftrightarrow}^{*}_{R}=\mathop{\leftrightarrow}^{*}_{E}$. We say that GTRS $R$ reaches a ground term $s\in T_{\Sigma}$ from a ground term $t\in T_{\Sigma}$ if $t\mathop{\rightarrow}_{R}^{*}s$ holds. We say that GTRS $R$ is terminating, (confluent, etc.) if $\mathop{\rightarrow}_{R}$ is terminating (confluent, etc.). A term $t\in T_{\Sigma}$ is irreducible for $R$ if it is irreducible for $\mathop{\rightarrow}_{R}$. If a term $t\in T_{\Sigma}$ has a unique $\mathop{\rightarrow}_{R}$ normal form, then the latter is denoted by $t{\downarrow}_{R}$. A GTRS $R$ is reduced if for every rewrite rule $l\dot{\rightarrow}r$ in $R$, $l$ is irreducible for $R-\{\,l\dot{\rightarrow}r\,\}$ and $r$ is irreducible for $R$. We recall the following important result.
Proposition 2.9 ([4, Theorem 2.14]).
Every reduced GTRS $R$ is convergent.
@@
Let $\Sigma$ be a signature, a bottom-up tree automaton (bta for short) over $\Sigma$ is a quadruple ${\cal A}=(\Sigma,A,R,A_{f})$, where $A$ is a finite set of states of arity 0, $\Sigma\cap A=\emptyset$, $A_{f}(\subseteq A)$ is the set of final states, $R$ is a finite set of rewrite rules of the following form:
$\sigma({b}_{1},\ldots,{b}_{k})\dot{\rightarrow}{b}$ with $k\in{\mathbb{N}}$, $\sigma\in\Sigma_{k}$, $b_{1},\ldots,b_{k},{b}\in A$.
We consider $R$ as a GTRS over $\Sigma\cup A$. The size of $\cal A$ is denoted by $\mathrm{size}({\cal A})$, and is $|A|+\mathrm{size}(R)$. The tree language recognized by $\cal A$ is $L({\cal A})=\{\,t\in T_{\Sigma}\mid(\exists{b}\in A_{f})\;t\mathop{\rightarrow}_{R}^{*}{b}\}$. We say that ${\cal A}=(\Sigma,A,R,A_{f})$ is a deterministic bta if for any $\sigma\in\Sigma_{k}$, $k\in{\mathbb{N}}$, ${b}_{1},\ldots,{b}_{k}\in A$, there is at most one rule with left-hand side $\sigma({b}_{1},\ldots,{b}_{k})$ in $R$.
@@
[7] Let $E$ and $F$ be GTESs over a signature $\Sigma$, and let $W=ST\langle E\cup F\rangle$. For every $p\in W$, we can construct in $O(n\,\mathrm{log}\,n)$ time the deterministic bottom-up tree automaton ${\cal A}=(\Sigma,C\langle E,F\rangle,R\langle E,F\rangle,\{\,p/_{\Theta\langle E,F\rangle}\,\})$, and $\cal A$} recognizes the tree language $p/_{\mathop{\leftrightarrow}^{*}_{E}}$.
@@
[10] Given two btas ${\cal A}=(\Sigma,A,R_{\cal A},A_{f})$ and ${\cal B}=(\Sigma,B,R_{\cal B},B_{f})$, we can decide whether $L({\cal A})\subseteq L({\cal B})$ in $O(\mathrm{size}({\cal A})\cdot\mathrm{size}({\cal B}))$ time.
@@
A directed pseudograph is a type of directed graph that allows for both parallel arcs (i.e. pairs of directed arcs with the same tail and the same head) and loops (i.e, directed arcs whose head and tail coincide) [19, top of page 4]. A directed pseudograph without parallel arcs (dpwpa for short) $G$ is an ordered pair $(V,A)$, where $V$ is a finite set, and $A\subseteq V\times V$ is a relation on $V$. The set $V$ is called the vertex set of $G$, and its elements are called vertices. The set $A$ is called the directed arc set of $G$, and its elements are called directed arcs. A walk in $G$ is an alternating sequence $w=x_{1}\,(x_{1},x_{2})\,x_{2}\,(x_{2},x_{3})\,x_{3}\,\ldots\,x_{k-1}\,(x_{k-1},x_{k})\,x_{k}$, $k\geq 1$, of vertices $x_{i}$ and arcs $(x_{j},x_{j+1})$. We say that $w$ is from vertex $x_{1}$ to vertex $x_{k}$. The length of $w$ is the number k1k-1 of its arcs [19, page 12]. A single vertex $x_{1}$ is considered as a walk from $x_{1}$ to $x_{1}$ of length 0. A trail is a walk $w$ in which all arcs are distinct. If the vertices of a trail $w$ are distinct, then $w$ is a path. If the vertices $x_{1},x_{2},\ldots,x_{k-1}$ of a trail $w$ are distinct, $k\geq 2$, and $x_{1}=x_{k}$, then $w$ is a cycle. We say that a vertex $x$ is reachable from a vertex $y$ if there exists a walk $w$ from vertex $y$ to vertex $x$. By definition, every vertex is reachable from itself. We say that a vertex $x$ is positive-step reachable from a vertex $y$ if there exists a walk $w$ of positive length from vertex $y$ to vertex $x$. Observe that for any vertices $x$ and $y$, the vertex $x$ is positive-step reachable from the vertex $y$ if and only if there exists a path of positive length or a cycle from the vertex $x$ to the vertex $y$.